40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7213     Entry time: Fri Aug 17 04:54:01 2012
Author: Yoichi, Koji 
Type: Summary 
Category: LSC 
Subject: PRMI Locking 

 To taste the strangeness of the current 40m PRC, I locked the PRMI with the guide of Koji.

We first aligned MICH by mostly tweaking ITMX, assuming that ITMY is in a good place as the Y-arm locks. MICH lock was stable.

Then we restored the IFO to the PRM_SBres mode. With a bit of alignment work on PRM and gain tweaking, the PRMI locked.

Yes, the beam spots look UGLY !

Also the PRMI was not so stable. Especially, when the alignment fluctuates, the optical gain changes and the loop becomes temporarily unstable. We took POP_DC as the guide for the gain change and normalized the PRCL error signal with it. To do this smoothly, we first changed the input matrix to route the PRCL error signal, which is REFL33_I, so that the signal also goes to the MC filter bank. Then with dtt, we monitored the spectra of the PRCL_IN1 and MC_IN1. We tweaked the value of the element in the normalization matrix for the MC path until the two spectra look the same (at this moment, the normalizing factor for the PRCL path was still zero). During this process, we noticed that the MC path signal (normalized by POP_DC) is noisier at above 500Hz. This was because the POP_DC has a large noise at high frequencies. So we put a low pass filter (100Hz 2nd order Butterworth) to the POP_DC filter bank to reduce the noise. Then the two spectra looked almost the same. The correct normalization factor found in this way was 0.03. So we put this number in the normalization matrix for PRCL. It did not break the PRMI lock.

 

After the normalization is turned on, the PRMI lock became somewhat more stable. However, the POP_DC was still fluctuating a lot, especially when the alignment is good. So I made a boost filter: 5Hz pole Q=2, 15Hz zero Q=1.5. I also made this filter automatically triggered when the PRMI is locked. This made the PRMI lock acquisition quicker. However, still the POP_DC fluctuation is large. It seems that the alignment of PRC is really fluctuating a lot.

 The current UGF of PRMI is about 150Hz with the phase margin over 50deg.

 

 

 

 

Attachment 1: AS_1029238601.jpg  9 kB  Uploaded Fri Aug 17 05:55:45 2012  | Hide | Hide all
AS_1029238601.jpg
Attachment 2: POP_1029238616.jpg  9 kB  Uploaded Fri Aug 17 05:55:55 2012  | Hide | Hide all
POP_1029238616.jpg
Attachment 3: REFL_1029238629.jpg  21 kB  Uploaded Fri Aug 17 05:56:03 2012  | Hide | Hide all
REFL_1029238629.jpg
Attachment 4: PRMI-OPLG.png  145 kB  Uploaded Fri Aug 17 06:17:26 2012  | Hide | Hide all
PRMI-OPLG.png
ELOG V3.1.3-